OpenStudy (anonymous):

A motorboat takes 4 hours to travel 128km going upstream. The return trip takes 2 hours going downstream. What is the rate of the boat in still water? What is the rate of the current?

OpenStudy (anonymous):

I was able to figure that one out. Now I've moved on to the following: two cyclists leave town 168 miles apart at the same time and travel toward each other. One cyclist travels 4 mi/h slower than the other. If they meet in 4 hours what is the rate of each cyclist?

OpenStudy (michele_laino):

If I call with v the speed of the motorboat with respect to the earth and with V the speed of the current, then the time nedded to go upstream is: \[\Large {t_a} = \frac{L}{{v - V}}\] whereas the time needed to got to downstream is: \[\Large {t_d} = \frac{L}{{v + V}}\]

OpenStudy (michele_laino):

where L= 128 Km

OpenStudy (michele_laino):

and ta= 4 hours, td= 2 hours

OpenStudy (anonymous):

Okay - I understand that...

OpenStudy (michele_laino):

so we can write this algebraic system: \[\Large \left\{ \begin{gathered} v - V = 32 \hfill \\ \hfill \\ v + V = 64 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

Okay

OpenStudy (michele_laino):

if we subtract the first equation from the second one, we get: \[\Large 2V = 32\] what is V?

OpenStudy (michele_laino):

@Juliette2120

OpenStudy (michele_laino):

here I call with v_A the speed of the first cyclist and with v_B the speed of the second cyclist, so I can write this: \[\Large {v_B} = {v_A} - 4\]

OpenStudy (michele_laino):

here is the situation described in your problem: